### v2

# Scissorhands: Scrub Data Influence via Connection Sensitivity in Networks

Abstract
Machine unlearning has become a pivotal task to erase the influence of data from a trained model. It adheres to recent data regulation standards and enhances the privacy and security of machine learning applications. In this work, we present a new machine unlearning approach Scissorhands. Initially, Scissorhands identifies the most pertinent parameters in the given model relative to the forgetting data via connection sensitivity. By reinitializing the most influential top-$k$ percent of these parameters, a trimmed model for erasing the influence of the forgetting data is obtained. Subsequently, Scissorhands fine-tunes the trimmed model with a gradient projection-based approach, seeking parameters that preserve information on the remaining data while discarding information related to the forgetting data. Our experimental results, conducted across image classification and image generation tasks, demonstrate that Scissorhands, showcases competitive performance when compared to existing methods.
Warning:
This paper contains explicit sexual imagery that may be offensive.
1 Introduction
In this work, we aim to propose an effective machine unlearning method. Under data regulations like the European Union General Data Protection Regulation (GDPR) [51] and the California Consumer Privacy Act (CCPA) [20], all users are granted the right to be forgotten. In machine learning, these legal provisions empower data owners with the right not only to withdraw their data from trained models but also to ensure that their data’s influence on these models is erased.
The most direct approach to accomplish this objective is to retrain the model from scratch, excluding the data requested for deletion from the training process. Retraining from scratch is typically considered the gold standard in the field of machine unlearning [50, 11]. Yet, this poses a challenge as numerous in-production models require prolonged training periods and substantial computing resources. While retraining is feasible, it is often impractical. Consequently, the proposal and development of efficient approximate unlearning methods [21, 49, 48, 12, 24] have become essential. Currently, most approximate unlearning techniques achieve forgetting by adding normally distributed noise to the parameters [17, 19, 18] or by estimating the influence of a particular data point on the model’s predictions [21, 36, 46, 35] based on the influence function method [10]. Jia et al. [25] demonstrate that model sparsity can help to unlearn and fuse model sparsity into the unlearning process.
Fan et al. [11] recently raised concerns regarding the instability in approximate unlearning methods, and highlighted that the current machine unlearning approaches, initially crafted for image classification tasks, fall short in effectively tackling the challenges of machine unlearning within the realm of image generation. The authors introduce a novel unlearning method SalUn that can effectively perform forgetting in both image classification and generation tasks. Their unlearning mechanism includes finding salient weights and then fine-tuning these weights using forgetting data assigned with random labels. However, studies like [56] show that the model can still learn the true data distribution even using data with random labels. While SalUn has shown state-of-the-art performance, it could potentially memorize information about forgetting data points, albeit associating them with random labels but still can be undesired in sensitive applications. In § 4, results also show that SalUn tends to memorize knowledge about the forgetting data.
This work. We present an unlearning approach, Scissorhands, designed to erase data influence in the classification models and eliminate a particular concept from a text-to-image model. To achieve this, inspired by [25], our key insight is to first erase the critical influence of forgetting data in models and then relearn. Unlike previous methods that employ the forgetting data with noisy labels as a part of relearning, we propose to use an efficient gradient projection method to relearn the critical features and patterns while ensuring the exclusion of influences associated with the forgetting data. Through a series of experiments and evaluations, including the classification on SVHN [37], CIFAR-10 and CIFAR-100 [28], CelebAMASK-HQ [29] datasets, as well as the open-source stable diffusion [42] text-to-image model, results demonstrate the viability and effectiveness of our technique in forgetting the influence of random samples, discrete classes, and sensitive content such as nudity.
2 Methodology
In this section, we propose Scissorhands, our unlearning framework that scrubs data from a model by eliminating their influence on model parameters. Throughout the paper, we denote scalars and vectors/matrices by lowercase and bold symbols, respectively (e.g., $a$, ${\bm{a}}$, and ${\bm{A}}$). We show a $d$-dimensional vector of ones by ${\bm{1}}_{d}$, and use $\odot$ to denote the Hadamard product.
2.1 Overview
Consider a model $f$ with parameters ${\bm{\theta}}\in{\mathbb{R}}^{d}$ trained on a dataset ${\mathcal{D}}=\{{\mathbf{x}}_{i},{\mathbf{y}}_{i}\}_{i=1}^{N}$. Suppose a subset of ${\mathcal{D}}$, denoted as ${\mathcal{D}}_{f}$, is requested for deletion. The remaining data is defined as ${\mathcal{D}}_{r}\coloneqq{\mathcal{D}}\setminus{\mathcal{D}}_{f}$. Our objective is to develop the unlearning algorithm to generate a scrubbed model $f_{u}$, effectively removing the influence of ${\mathcal{D}}_{f}$ from model $f$, while maintaining its utility over ${\mathcal{D}}_{r}$.
In Scissorhands, as shown in Figure 1, we initially identify critical parameters w.r.t. the forgetting data via connection sensitivity analysis [30, 31]. We then reinitialize the top-$k$ percent of these key parameters, resulting in a trimmed model with parameters ${\bm{\theta}}_{t}$. This step aims to diminish the model’s memory of the forgetting data ${\mathcal{D}}_{f}$. Nevertheless, this process runs the risk of erasing valuable information from the data set ${\mathcal{D}}_{r}$ that we aim to retain.
To address this, the next phase of Scissorhands concentrates on restoring the performance of the trimmed model on remaining data ${\mathcal{D}}_{r}$, while ensuring that the influence of the forgetting data ${\mathcal{D}}_{f}$ remains excluded. We accomplish this balance and obtain the final scrubbed model with parameters ${\bm{\theta}}^{\star}$ through a gradient projection-based approach.
Figure 1: Overview of Scissorhands. Given a model with parameter ${\bm{\theta}}$ trained using ${\mathcal{D}}$, in the trimming process, key parameters w.r.t. the forgetting data ${\mathcal{D}}_{f}$ would be pinpointed via Equation 2, and top-$k$of these parameters would be re-initialized for erasing the influence of ${\mathcal{D}}_{f}$, resulting in the trimmed model. In the repairing process, a gradient projection-based method is adopted for restoring the trimmed model utility over the remaining data ${\mathcal{D}}_{r}$ while excluding information associated with ${\mathcal{D}}_{f}$.
2.2 Trimming
To effectively identify salient parameters of a network with respect to ${\mathcal{D}}_{f}$, one has to establish a good criterion to determine such salient connections. A commonly used criterion is the magnitude of the weights, with those exceeding a certain threshold deemed salient as suggested in [23]. Our requirement is more nuanced, as we need a measure that specifically discerns saliency based on ${\mathcal{D}}_{f}$. Therefore, we adopt a single-shot approach as proposed in [30], which we outline briefly to ensure clarity and completeness in our methodology. Consider
$\texttt{s}_{j}({\mathcal{D}})\coloneqq\mathbb{E}_{{\bm{x}},{\bm{y}}\sim{\mathcal{D}}}\Big{[}\ell({\bm{\theta}};{\bm{x}},{\bm{y}})-\ell(({\bm{1}}_{d}-{\bm{e}}_{j})\odot{\bm{\theta}};{\bm{x}},{\bm{y}})\Big{]},$ (1)
which measures the influence of parameter $j\in\{1,\ldots,d\}$ on a model in terms of the empirical risk for a given dataset ${\mathcal{D}}$. Here, $\ell:\mathbb{R}^{d}\times\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+}$ is the loss of the model and ${\bm{e}}_{j}$ is the indicator vector of element $j$ (i.e., a binary vector with its $j$-th component set to one). Note that computing $\texttt{s}_{j}$ for each $j\in\{1,\ldots,d\}$ is prohibitively expensive, as it demands $d+1$ forward passes through the dataset. An approximation to Equation 1 [23] is in the form of
$\texttt{s}_{j}({\mathcal{D}})\approx\mathbb{E}_{{\bm{x}},{\bm{y}}\sim{\mathcal{D}}}\Big{[}\frac{\partial\ell({\bm{\theta}};{\bm{x}},{\bm{y}})}{\partial{\bm{\theta}}_{j}}{\bm{\theta}}_{j}\Big{]}\;.$ (2)
Detailed proofs can be found in Section 7. Equation 2 defines the sensitivity of parameter $j$ as the average (i.e., expectation) of a product, the gradient of the loss and the current value of the parameter. As such, $\texttt{s}_{j}$ encodes the loss sensitivity according to the parameter’s magnitude. Intuitively, a parameter with a large value that significantly affects the loss, as captured by the gradient, is considered more influential or salient. The form in Equation 2 offers several benefits, most importantly, the ability to calculate the sensitivity for all parameters with just a single sweep through the dataset.
To scrub the influence of the forgetting data ${\mathcal{D}}_{f}$ in the model, we first obtain $\texttt{s}_{j}({\mathcal{D}}_{f})$, saliency of parameters w.r.t. the forgetting data. We then re-initialize the top-$k\%$ of the parameters based on their saliency rankings. This process is akin to performing a targeted ‘lobotomy’ on the model specifically concerning the data ${\mathcal{D}}_{f}$, thereby selectively erasing their influence. Unfortunately, this aggressive approach of reinitializing parameters can detrimentally impact the model’s performance on the remaining dataset ${\mathcal{D}}_{r}$, which we will correct in the next phase of Scissorhands.
Remark 1.
In practice, we have observed that even a small subset of ${\mathcal{D}}_{f}$ can be sufficient for the trimming. For example, to unlearn a large set of data on CIFAR-100, our algorithm outperforms baselines by utilizing approximately 3% of the forgetting data for trimming.
Remark 2.
Initializing weights of the trimmed neurons can take various forms. Empirically, we observed initializing with a uniform distribution is particularly effective. Details on the influence of different initialization strategies and the choice of ktrimming on the model’s performance will be discussed in § 4.2.
2.3 Repairing
Following the parameter-trimming process, aimed at mitigating the influence of the forgetting data ${\mathcal{D}}_{f}$, a challenge presents itself: the potential erasure of crucial information associated with the remaining data ${\mathcal{D}}_{r}$. A straightforward solution to recover the model utility is to relearn the model using the remaining data. However, this approach risks biasing the model towards features of the remaining data and inadvertently reintegrating information about the forgetting samples. In the quest to ensure that models retain essential information from the remaining data ${\mathcal{D}}_{r}$ while effectively forgetting ${\mathcal{D}}_{f}$, our strategy involves maximizing the loss over the forgetting data ${\mathcal{D}}_{f}$ while concurrently minimizing the loss over the remaining data ${\mathcal{D}}_{r}$. Therefore, to achieve the balance between these goals, we propose the objective of an efficient practice for unlearning:
$L({\bm{\theta}},{\mathcal{D}}_{r},{\mathcal{D}}_{f})\coloneqq\mathcal{L}({\bm{\theta}};{\mathcal{D}}_{r})-\lambda\mathcal{L}({\bm{\theta}};{\mathcal{D}}_{f})\;.$ (3)
We aim to optimize Equation 3 w.r.t. ${\bm{\theta}}$ through gradient descent, ensuring minimal reintroduction of information about ${\mathcal{D}}_{f}$. This is particularly challenging when there are similarities between samples in ${\mathcal{D}}_{f}$ and ${\mathcal{D}}_{r}$. To achieve this desideratum, it is vital that updates to the model do not improve (i.e., reduce) $\mathcal{L}({\bm{\theta}};{\mathcal{D}}_{f})$. To this end, we define ${\bm{g}}_{f}=\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\theta}};{\mathcal{D}}_{f})$ as the gradient direction for ${\mathcal{D}}_{f}$ and ${\bm{g}}_{o}=\nabla_{{\bm{\theta}}}L({\bm{\theta}},{\mathcal{D}}_{r},{\mathcal{D}}_{f})$ as the gradient direction for optimizing Equation 3. The optimal descent direction ${\bm{g}}\in\mathbb{R}^{d}$ should thus satisfy two criteria: 1. It should exhibit maximum similarity to ${\bm{g}}_{o}$ to ensure swift convergence. The notion of similarity can be captured by ${\bm{g}}_{o}$ using $\left\|{\bm{g}}-{{\bm{g}}_{o}}\right\|^{2}_{2}$. 2. It should not align with ${\bm{g}}_{f}$ to prevent improving $\mathcal{L}({\bm{\theta}};{\mathcal{D}}_{f})$. This can be mathematically formalized as $\langle{\bm{g}}_{f},{\bm{g}}\rangle\leq 0$. Consequently, after computing ${\bm{g}}_{o}=\nabla_{{\bm{\theta}}}L({\bm{\theta}},{\mathcal{D}}_{r},{\mathcal{D}}_{f})$ and ${\bm{g}}_{f}=\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\theta}};{\mathcal{D}}_{f})$ in each iteration, we formulate the following optimization challenge to identify the model update direction:
$\operatorname*{arg\,min}_{{{\bm{g}}}}\quad$ $\frac{1}{2}\left\|{\bm{g}}-{{\bm{g}}_{o}}\right\|^{2}_{2},$
s.t. $\langle{\bm{g}}_{f},{\bm{g}}\rangle\leq 0.$ (4)
This problem may be addressed using the Frank-Wolfe algorithm. Given the high dimensionality of gradients in neural networks, directly addressing the constraint optimization in Section 2.3 could become overwhelming. Inspired by [34], we propose to make use of the dual formulation of the problem Section 2.3 as:
$\sup_{v}\quad$ $\frac{1}{2}{\bm{g}}_{f}^{\top}{\bm{g}}_{f}v^{2}-{\bm{g}}_{o}^{\top}{\bm{g}}_{f}v,$
s.t. $v\geq 0.$ (5)
Then, the optimal descent direction ${\bm{g}}$ is given from the solution $v^{*}$ in Section 2.3 as ${\bm{g}}={\bm{g}}_{o}-v^{*}{\bm{g}}_{f}$. Detailed proofs can be found in  Section 7. By doing so, we ensure a balance between unlearning and retaining utility.
Remark 3.
The primal problem (cf. Section 2.3) involves optimizing over the high-dimensional space of gradient vectors, while the dual problem (cf. Section 2.3) simplifies this problem to optimize over a single scalar variable $v\in\mathbb{R}$. This dimensionality reduction can significantly decrease the computational complexity, especially for large-scale neural networks.
2.4 Algorithm Description
Algorithm 1 The procedure of Scissorhands.
1:procedure Trimming(${\bm{\theta}},{\mathcal{D}}_{f}$)
2:     ${\bm{X}},{\bm{y}}\leftarrow\{{\mathbf{x}}_{i},{\mathbf{y}}_{i}\sim{\mathcal{D}}_{f}|i\in[B]\}$.
3:     // Compute connection sensitivity
4:     Get $\texttt{s}_{j}({\bm{\theta}};{\bm{X}},{\bm{y}}),\hskip 1.99997pt\forall j\in[1,d]$ using Eq. 2.
5:     // Re-initialization
6:     Re-initialize parameters according to the
7:       top-$k$value of $\texttt{s}_{j}$.
8:     return ${\bm{\theta}}_{t}$
9:end procedure
10:
11:procedure Repairing(${\bm{\theta}}_{t},{\mathcal{D}}_{f},{\mathcal{D}}_{r}$)
12:     ${\bm{\theta}}^{0}={\bm{\theta}}_{t}$.
13:     for $e=0$ to $E-1$ do
14:         Get $L({\bm{\theta}},{\mathcal{D}}_{r},{\mathcal{D}}_{f})$ (cf. Eq. 3).
15:         Get ${\bm{g}}_{o}$ and ${\bm{g}}_{f}$.
16:         // Get the optimal direction
17:         if $\langle{\bm{g}}_{o},{\bm{g}}_{f}\rangle>0$ then
18:              Compute $v^{*}$ (cf. Sec. 2.3).
19:              ${\bm{g}}={\bm{g}}_{o}-v^{*}{\bm{g}}_{f}$.
20:              ${\bm{g}}_{o}\leftarrow{\bm{g}}$.
21:         end if
22:         ${\bm{\theta}}^{e+1}\leftarrow{\bm{\theta}}^{e}-\eta{\bm{g}}_{o}$.
23:     end for
24:     return ${\bm{\theta}}^{*}={\bm{\theta}}^{E}$.
25:end procedure
Algorithm 1 describes the procedure of Scissorhands in detail. We first identify the important parameters w.r.t. the forgetting data in single-shot via connection sensitivity in networks, then re-initialize these parameters for erasing the influence of the forgetting data ${\mathcal{D}}_{f}$ in the given model. To repair the performance of the trimmed model on the remaining data ${\mathcal{D}}_{r}$, while ensuring the exclusion of information associated with the forgetting data ${\mathcal{D}}_{f}$, we employ the gradient projection-based optimization algorithm. By doing so, we can scrub the influence of the forgetting data while retaining model utility on the remaining data.
3 Related Work
The development of efficient machine unlearning methods [43, 26, 3, 16, 2, 40, 55, 21, 17, 35, 46, 25, 48, 38] has gained prominence. Applications of machine unlearning span various domains, including regression tasks [48], federated learning [33, 32, 22, 54, 52], graph neural networks [4, 7], and diffusion models [14, 57, 24, 15, 11], as well as scenarios where training data are not available [49, 8].
Retraining the model from scratch without forgetting data is typically considered the standard gold unlearning algorithm [50, 11]. However, this is often deemed impractical as most in-production models require extensive training periods and considerable computing resources. While fine-tuning the model for a new task might lead to forgetting previous knowledge (i.e., catastrophic forgetting) [34], it fails to adequately erase the data influence in the models. Approximate unlearning methods, as such, become attractive alternatives. We will briefly discuss these methods in image classification and image generation.
Unlearning in Classification Models. Most unlearning algorithms are based on the influence function [46, 21, 36] and the Fisher Information Matrix (FIM) [17, 18, 19]. Influence function-based methods estimate the influence of the particular training data points on the model’s predictions, and Fisher unlearning methods assume that the unlearned model and the retrained model are close to each other, Golatkar et al. [17] leverage FIM and hide the difference between the unlearned models and retrained models via adding noise. Golatkar et al. [17] also present an upper bound for the amount of retained information, offering a quantifiable measure of the effectiveness of the unlearning process. Such approximate unlearning methods need to compute the Hessian matrix, or FIM w.r.t. the data. This may render them impractical for certain scenarios, such as federated learning, where computational resources are distributed and limited. To mitigate these computational demands, Mehta et al. [35] introduce L-CODEC, a strategy aimed at pre-selecting a subset of parameters, allowing for more efficient computation. Jia et al. [25] recently propose to fuse the model sparsity into the unlearning algorithm, helping improve approximate unlearning methods, and achieving effective and efficient unlearning effects.
Concept Erasure in Generative Models. The advent of generative models, particularly those converting text to images, has been a significant milestone in the field of artificial intelligence. A notable concern is the risk of it being tainted or manipulated [6, 41], resulting in not-safe-for-work (NSFW) generations [45], as these models leverage training data from a wide array of open sources. To address these challenges, data censoring [13, 1, 39, 44] to exclude black-listed images is employed. Studies [14, 45] introduce methods to update the models away from the inappropriate concepts. Heng and Soh [24] recently propose to adopt Elastic Weight Consolidation (EWC) and Generative Replay (GR) to effectively unlearn without access to the training data for a wide range of generative models.
Unlearning across Domains. Recently, Fan et al. [11] highlight the difficulty in cross-domain applicability of the machine unlearning algorithms, i.e., existing machine unlearning methods designed for classification tasks cannot be effective when applied to image generation. To address this gap, they introduce a novel method, Saliency unlearning (SalUn), which shifts attention from the entire model to target parameters. SalUn achieves state-of-the-art performance and is effective in both image classification and image generation tasks.
In this work, we propose Scissorhands which could achieve improved performance compared to existing methods across various scenarios. Inspired by [25] that apply $\ell_{1}$ regularization to simply fuse sparsity into the unlearning methods, we identify and re-initialize the key model parameters w.r.t. the forgetting data via connection sensitivity. Concurrently, Fan et al. [11] also explore unlearning via salience scores, determining weight salience via the gradient of the forgetting loss w.r.t. model parameters, and employing random labeling for model fine-tuning. While there are similarities in our overarching goals, the methodologies present unique perspectives and solutions to the problem of unlearning. Fan et al. [11] only focus on the salient weights, while our mechanism aims to first ‘destroy’ the key information and then relearn for the whole model. Additionally, using the forgetting data with random labels could potentially lead to memorizing information about the forgetting data points [56]. In contrast, Scissorhands excludes knowledge about the forgetting data via a gradient projection-based approach, resulting in the effective erasure of information about the forgetting data while preserving the knowledge about the retained data.
4 Experimental Evaluation
In this section, we illustrate how Scissorhands effectively eliminates the data influence in the models. For sample-wise forgetting, where the forgetting data shares the same distribution as the training data, we evaluate Scissorhands on SVHN [37], CIFAR-10 and CIFAR-100 [28]. We further extend our evaluation on CelebAMask-HQ [29] where targeting to remove the entirety of specific identities. Additionally, our experimentation encompasses the open-source text-to-image model, Stable Diffusion v1.4 [42], which is conditioned on CLIP text embeddings through the cross-attention mechanism. Further detailed experimental setups and additional results can be found in Sec. 8.
Baselines.
We primarily compare against the following standard baselines that are frequently employed in machine unlearning, alongside recently proposed SOTA methods: (i) Retrain: models obtained by only retraining from scratch on ${\mathcal{D}}_{r}$. This will provide the performance of the oracle for us. (ii) Fine-tuning (FT) [53]: models that are fine-tuned on ${\mathcal{D}}_{r}$, i.e., taking advantage from catastrophic forgetting in neural networks to unlearn. (iii) Gradient ascent (GA) [50]: gradient ascent on ${\mathcal{D}}_{f}$. This will provide a simple way of unlearning by making the performance of the model worse on ${\mathcal{D}}_{f}$. (iv) Influence unlearning (IU) [27]: utilizes the influence function method to estimate the change in model parameters when transitioning from the unscrubbed model to the retrained model. (v) Boundary shrink (BS) [5] and (vi) Boundary expanding (BE) [5]: shift the decision boundary of the original model to imitate that of the retrained model. (vii) $\ell_{1}$-sparse [25]: fine-tuning models on ${\mathcal{D}}_{r}$ with $\ell_{1}$-norm sparse regularization. (viii) Saliency unlearning (SalUn) [11]: adopt weight saliency and random labeling for unlearning.
Metrics.
To assess the unlearning algorithms, we employ several metrics: (i) Accuracy: the accuracy of the model on ${\mathcal{D}}_{f}$ (denoted as $\text{Acc}_{{\mathcal{D}}_{f}}$), ${\mathcal{D}}_{r}$ (denoted as $\text{Acc}_{{\mathcal{D}}_{r}}$) and the test set (denoted as $\text{Acc}_{{\mathcal{D}}_{t}}$). It provides insight into how well the model performs after undergoing the unlearning process. (ii) Membership inference attack (MIA): a standard metric for verifying the unlearning effect. A classifier $\phi$ is trained using both training data (marked with a label of 1, indicating data points included in training) and test data (marked with a label of 0, indicating data points not seen during training), and subsequently assessed on ${\mathcal{D}}_{f}$. An effective unlearning method should make it challenging to infer whether a particular sample was part of the training set, mitigating potential privacy breaches. In this context, we define MIA as $P_{\phi}({\bm{y}}=0|{\bm{x}}_{f})$, representing the probability assigned by the classifier $\phi$ that a given sample ${\bm{x}}_{f}$ was not included in the training data. (iii)Avg. Gap [11]: the average of the performance gaps measured in accuracy-related metrics: Avg. Gap $=(|\text{Acc}_{{\mathcal{D}}_{t}}-\text{Acc}^{\ast}_{{\mathcal{D}}_{t}}|+|\text{Acc}_{{\mathcal{D}}_{f}}-\text{Acc}^{\ast}_{{\mathcal{D}}_{f}}|+|\text{Acc}_{{\mathcal{D}}_{r}}-\text{Acc}^{\ast}_{{\mathcal{D}}_{r}}|+|\text{MIA}-\text{MIA}^{\ast}|)/4$. where $\text{Acc}^{\ast}_{{\mathcal{D}}_{t}},\text{Acc}^{\ast}_{{\mathcal{D}}_{f}},\text{Acc}^{\ast}_{{\mathcal{D}}_{r}}$ and $\text{MIA}^{\ast}$ are metric values of the retrained model. A better performance of an unlearning method corresponds to a lower performance gap with retraining, measuring how close the unlearned model is to the retrained model. (iv) Run-time efficiency (RTE) [25]: evaluate the computational efficiency of the unlearning method. Specifically, we have $\text{RTE}={\text{T}}/{\text{T}_{r}}$, where T and $\text{T}_{r}$ present the time for the unlearning methods to get the scrubbed models and the time for retraining from scratch with ${\mathcal{D}}_{r}$, respectively. An efficient unlearning method should aim for minimal computational overhead compared to the baseline. (v) Relearn time [17, 8, 48]: the number of epochs to fine-tune the scrubbed models for regaining performance on ${\mathcal{D}}_{f}$ (i.e., reach the original models’ accuracy on ${\mathcal{D}}_{f}$). The larger the number, the more effective the erasure.
4.1 Results on Classification Task
Table 1: Quantitative results for forgetting 10% data on the SVHN, CIFAR-10 and CIFAR-100 datasets.
Method $\text{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\text{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ $\text{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ MIA()($\uparrow$)( ↑ ) Avg. Gap
CIFAR-100 Retrain 75.13$\pm$0.85 74.69$\pm$0.08 99.98$\pm$0.01 50.22$\pm$0.62 -
[…]
GA [50] 98.00$\pm$1.34 75.59$\pm$0.11 98.24$\pm$1.16 5.00$\pm$2.25 17.68
IU [27] 95.67$\pm$4.82 72.13$\pm$4.58 96.14$\pm$4.51 9.43$\pm$5.98 16.93
BE [5] 97.94$\pm$1.38 74.16$\pm$0.09 98.12$\pm$1.24 7.60$\pm$3.05 16.96
BS [5] 97.65$\pm$1.48 73.20$\pm$0.18 97.93$\pm$1.30 8.24$\pm$3.23 17.01
$\ell_{1}$-sparse [25] 96.35$\pm$0.67 70.06$\pm$0.46 96.35$\pm$0.67 21.33$\pm$1.95 14.59
SalUn [11] 88.56$\pm$1.18 71.34$\pm$0.48 99.40$\pm$0.35 74.66$\pm$2.48 10.45
Ours 68.76$\pm$1.81 73.17$\pm$0.24 99.24$\pm$0.30 42.42$\pm$2.06 4.11
CIFAR-10 Retrain 94.81$\pm$0.53 94.26$\pm$0.14 100.0$\pm$0.00 13.05$\pm$0.64 -
[…]
GA [50] 99.66$\pm$0.23 94.57$\pm$0.01 99.62$\pm$0.25 0.91$\pm$0.29 4.42
IU [27] 98.08$\pm$2.10 91.91$\pm$2.73 98.01$\pm$2.26 4.01$\pm$3.44 4.16
BE [5] 99.41$\pm$0.38 93.79$\pm$0.15 99.41$\pm$0.38 16.16$\pm$0.78 2.19
BS [5] 99.60$\pm$0.25 94.24$\pm$0.07 99.56$\pm$0.54 4.46$\pm$0.33 3.46
$\ell_{1}$-sparse [25] 94.17$\pm$0.49 90.64$\pm$0.52 96.64$\pm$0.54 11.87$\pm$0.61 2.20
SalUn [11] 98.07$\pm$0.42 93.92$\pm$0.25 99.89$\pm$0.07 17.93$\pm$0.37 2.15
Ours 95.40$\pm$1.48 92.92$\pm$0.48 98.93$\pm$0.57 9.56$\pm$2.13 1.62
SVHN Retrain 91.81$\pm$1.11 91.17$\pm$1.77 97.73$\pm$1.12 15.74$\pm$1.28 -
[…]
GA [50] 98.41$\pm$0.23 92.87$\pm$0.06 98.52$\pm$0.29 5.96$\pm$0.38 4.72
IU [27] 92.47$\pm$1.62 86.92$\pm$2.06 93.36$\pm$1.82 17.00$\pm$2.67 2.63
BE [5] 98.48$\pm$0.33 92.62$\pm$0.05 98.44$\pm$0.29 6.77$\pm$0.36 4.45
BS [5] 98.29$\pm$0.32 92.48$\pm$0.02 98.36$\pm$0.29 6.74$\pm$0.39 4.36
$\ell_{1}$-sparse [25] 97.79$\pm$0.10 93.59$\pm$3.18 99.49$\pm$0.23 7.57$\pm$0.29 4.65
SalUn [11] 96.29$\pm$4.80 93.59$\pm$3.18 97.41$\pm$4.96 24.47$\pm$4.08 3.99
Ours 91.62$\pm$1.18 93.05$\pm$0.56 99.51$\pm$0.14 17.62$\pm$2.54 1.43
Table 2: Quantitative results for forgetting 10% identities on the CelebAMask-HQ.
Method $\text{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\text{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ $\text{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ MIA()($\uparrow$)( ↑ ) Avg. Gap
Retrain 0.00$\pm$0.00 87.02$\pm$0.80 99.96$\pm$0.01 100.$0$
[…]
GA [50] 87.60$\pm$8.71 81.22$\pm$2.11 99.74$\pm$0.26 51.37$\pm$5.96 35.56
IU [27] 88.92$\pm$10.25 70.24$\pm$11.77 95.27$\pm$5.07 29.59$\pm$18.59 45.20
BE [5] 69.07$\pm$2.73 44.11$\pm$2.08 95.58$\pm$1.23 46.24$\pm$5.90 42.53
BS [5] 98.18$\pm$1.92 81.92$\pm$0.27 99.86$\pm$0.03 45.93$\pm$5.11 39.36
$\ell_{1}$-sparse [25] 98.81$\pm$0.72 89.37$\pm$0.70 99.97$\pm$0.00 76.78$\pm$5.66 31.10
SalUn [11] 0.00$\pm$0.00 78.36$\pm$1.34 96.90$\pm$1.11 100.0$\pm$0.00 2.93
Ours 1.52$\pm$2.73 80.18$\pm$6.60 97.20$\pm$3.81 99.83$\pm$0.35 2.82
Figure 2: Visualizations of regions where models focus on generated by GradCAM [47]. Best viewed in color.
In this experiment on SVHN, CIFAR-10, and CIFAR-100, we try to forget randomly selected 10% of the training data; on CelebAMask-HQ, we attempt to forget randomly selected 10% identities among 307 identities. We further apply GradCAM [47] to visualize regions where models focus on w/ and w/o machine unlearning algorithms. In brief, the results suggest that Scissorhands has successfully induced forgetting for the relevant samples and classes, with minor degradation in model performance over the remaining data and classes. We will discuss the results in depth below.